OpenStudy (kirbykirby):

How can I solve this system of equations for any \(n\): \[1.25a_1+0.5a_2=0.5\\ 0.5a_1+1.25a_2+0.5a_3=0\\ 0.5a_2+1.25a_3+0.5a_4=0\\ 0.5a_3+1.25a_4+0.5a_5=0\\ ...\\ 0.5a_{n-2}+1.25a_{n-1}+0.5a_n=0\\ 0.5a_{n-1}+1.25a_n=0\]

OpenStudy (kirbykirby):

The best I could do is get "some sort" of recursive formula: \[a_1=0.4-0.4a_2\\ a_2=1-2.5a_1\\ a_k=-a_{k-2}-2.5a_{k-1} \text{ for } 3 \le k \le n-1 \\ a_n=-0.4a_{n-1}\\ \] But doesn't help too much since a1 depends on a2, and a2 depends on a1 :(

OpenStudy (perl):

interesting

ganeshie8 (ganeshie8):

n equations and n unknowns, so solvable i may try something like below :- 1) eliminate \(a_1\) from first two equations : (2) - (1)/2.5 ----------(I) 2) eliminate \(a_2\) from above resultant equation and third equation : (3) - (I) --------(II) 3) keep going, final two equations will have two unknowns : \(a_{n-1}\) and \(a_n\)

ganeshie8 (ganeshie8):

you're looking for deriving a formula for finding last two equations is it ?

OpenStudy (kirbykirby):

well basically some sort of equation(s) that will provide a solution to \(a_1\) to \(a_n\)

ganeshie8 (ganeshie8):

looks like bit involved task...

OpenStudy (kirbykirby):

Yeah :( I'm not sure how to do it exactly. and i can't really use software since I have to find it for any general n

OpenStudy (kirbykirby):

or can I? o_o I'm sorry I am so brain-dead. I have been up all night ;_;

ganeshie8 (ganeshie8):

another thing to try is (painful too) :- 1) setup a nxn matrix (u wil get a symmetric matrix here) 2) row reduce to triangular 3) setup equations

ganeshie8 (ganeshie8):

is this question from linear algebra... or just a general one ?

OpenStudy (loser66):

for this \(0.5 a_{n-2} + 1.25a_{n-1}+0.5a_n=0\) you can write down back ward and make it have a integer coefficient by time all by 4, \[2a_{n +2}+ 5a_{n+1} +2a_n=0\] characteristic equation is \(2r^2+5r +2=0\) solve for it, you have 2 roots \(r_1= -0.5\) and \(r_2 = -2\) so \(a_n = C_0(-0.5)^n + C_1(-2)^n\)

OpenStudy (loser66):

If you have 2 initial conditions for \(a_0\) and \(a_1\), you can solve for \(C_0\) and \(C_1\) If you don't have , just stop there. You are OK. Is this from discrete math?

OpenStudy (kirbykirby):

actually this is a sub result I am getting from my stats problem

OpenStudy (kirbykirby):

I assume not many people here have done upper-year stats, but it's a forecasting problem and I need to solve these equations as I progress through the problem.

OpenStudy (kirbykirby):

the thing is I don't have any initial conditions so this was giving me a lot of grief

OpenStudy (amistre64):

the initial conditions should result in arbitrary constants i believe

OpenStudy (amistre64):

or at least one arbitrary constant with which to define when a specific initial condition is known

OpenStudy (kirbykirby):

thanks everyone for their responses!

OpenStudy (loser66):

hope this helps

OpenStudy (kirbykirby):

Oh thanks very much^!